OpenStudy (anonymous):

Simple question?√(2x+13)-5=x (Just need to check my answer)

OpenStudy (anonymous):

Ok I know the first step is to add 5 to both sides: √(2x+13)=5x. But I can't remember how to go from there.

OpenStudy (mathstudent55):

Correct. First add 5 to both sides. Then, square both sides to get rid of radical. Then solve equation. Then check all answers in original equation because squaring both sides may introduce extraneous solutions.

OpenStudy (mathstudent55):

BTW, right side is x + 5, not 5x

OpenStudy (anonymous):

That's what I figured but I wasn't sure if it should look like this after it as squared: 4x+169=25+x. (Btw thanks for the correction :) )

OpenStudy (mathstudent55):

Not quite. When you square a binomial there is another term. (a + b)^2 = a^2 + 2ab + b^2

OpenStudy (mathstudent55):

The left side is in a square root, so squaring it just removes the square root.

OpenStudy (mathstudent55):